An Ice cube at -35o C is added to 250 g of Apple juice in a 100 g glass cup both
at 25o C, Find the mass of the ice that make the hole system at 5o C where the
specific heat capacity for apple juice is 3.64 kJ/kg. o C, for the water is 4.18
kJ/kg. o C and for glass is 0.2 kJ/kg. o C while for ice is 2.108 kJ/kg. o C and
the fusion specific latent heat is 3.36 kJ/kg. o C.
1
Expert's answer
2020-07-27T09:34:50-0400

Let us calculate the total amount of thermal energy and write the equation of balance:

"c_im_i\cdot35K + L_im_i +c_wm_i\cdot 5K = c_jm_j\cdot 20K + c_gm_g\cdot 20K
\\

\\"


We can see that we have an equation on mi. Let us substitute all numerical values and get the mass of ice.

We obtain the value of mass equal to 43 gramms.